A jar contains 15 blue and 10 red marbles..
a) If one marble is drawn at random, what is the probability that it is red?
b) If two marbles are drawn randomly, what is the probability that the first is red and the second blue if the first marble is replaced in the jar prior to drawing the second marble?
c) What is the probability that both are red if the first marble is not returned to the jar prior to drawing the second marble?
: A jar contains 15 blue and 10 red marbles..
a) If one marble is drawn at random, what is the probability that it is red?
b) If two marbles are drawn randomly, what is the probability that the first is red and the second blue if the first marble is replaced in the jar prior to drawing the second marble?
c) What is the probability that both are red if the first marble is not returned to the jar prior to drawing the second marble?

Answer by stanbon(18984)

A jar contains 15 blue and 10 red marbles..
a) If one marble is drawn at random, what is the probability that it is red?
There are 10 ways to succeed; there are 25 possible outcomes
P(red) = 10/25 = 2/5
-----------------------------
b) If two marbles are drawn randomly, what is the probability that the first is red and the second blue if the first marble is replaced in the jar prior to drawing the second marble?
If the 1st marble is replaced the selections are independent.
If events A and B are independent, P(A and B) is P(A)*P(B)
P(red and blue with replacement) = P(red)*P(blue) = 10/25 * 15/25 = 150/626
= 75/313
----------------------------
c) What is the probability that both are red if the first marble is not returned to the jar prior to drawing the second marble?
If the 1st marble is not replaced the selections are dependent.
If events A and B are dependent, P(A and B) = P(A)*P(B |A)
P(red and red without replacement) = P(red)*P(red | red) = (10/25)*(9/24)
= 3/20
===================
Cheers,
Stan H.